Find the equation of the line in standard form through the points (-3, -1) and (-5,7).
Answer by mananth(16946)

x1 y1 x2 y2
-3 -1 -5 7

slope m = (y2-y1)/(x2-x1)
(7-(-1) )/(-5-(-3))
(8/-2)
m=-4.00

Plug value of the slope and point (-3,-1) in
Y = m x + b
-1.00 = 12 + b
b= -1 12
b= -13
So the equation will be
Y = -4 x -13
4x+y=-13
m.ananth@hotmail.ca